OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (-1, 6) and has a slope of -3.

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ -1}}\quad ,&{\color{blue}{ 6}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= -3 \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (anonymous):

y - 6 = -3x - 3 y - 6 = -3(x + 1) y = -3x + 3 3x + y = 3

OpenStudy (anonymous):

y-6=m(x-(-1)

OpenStudy (jdoe0001):

yeap, recall that the slope = m = -3

OpenStudy (jdoe0001):

thus y-6 = -3(x-(-1) y -6 = -3(x+1)

OpenStudy (anonymous):

ok whats next

OpenStudy (anonymous):

oh nevermind lol